A cell’s immune response to an infection depends on a sequence of chemical mes-
sengers. In researching how such chemical messengers are transported from outside into the cell,

Rutger D. Luteijn and Shivam A. Zaver et al. measured how folic acid (vitamin B9) and other med-
ications inhibit the cell’s uptake (absorption) of immunoreactive chemical messengers [2]. Table 2

summarizes some of their data.© 2020 Y. Nievergelt MATH 241-001 Problems 1 AND 2 due by midnight PDT on 30 July 2 Example 2 A cells immune response to an
Answer #1

(2.1) We just evaluate the formula for U at \beta = 4 ,

U(4) =\sum_{k=1}^4 (y_k - 4 t_k)^2
= \left( \frac{266.9}{3} -4\times 30 \right )^2 + \left( \frac{705.1}{3}-4\times 60 \right )^2+ \left(\frac{1772.5}{3} - 4\times 150 \right )^2+ \left( \frac{4165.5}{5}-4\times 300\right )^2
= 36\,604 .

(2.2) We have

U'(\beta) = \left( \sum_{k=1}^4 (y_k - \beta t_k)^2 \right )'
= \sum_{k=1}^4 \left((y_k - \beta t_k)^2 \right )'
= \sum_{k=1}^4 2\times (y_k - \beta t_k)\times (-t_k)
=- 2 \sum_{k=1}^4 t_k (y_k - \beta t_k)

and therefore

U'(4) = -2 \sum_{k=1}^4 t_k (y_k - 4 t_k) = -10\,789.

(2.3) U(\beta) is minimized when U'(\beta) = 0 ,

0= U'(\beta)
=- 2 \sum_{k=1}^4 t_k (y_k - \beta t_k)

We can cancel the -2 factor and get

0 = \sum_{k=1}^4 t_k (y_k -\beta t_k)
0 = \sum_{k=1}^4 t_k y_k - \beta\sum_{k=1}^4 t_k^2

\beta = \frac{\sum y_k t_k}{\sum t_k^2} = 4.4611.

(2.4) The value \beta = 4.4611 corresponds to the slope \tilde \beta_y of the line passing thought the origin that fits the data the best according to least squares.

(2.5) Here is the plot of U vs \beta

X106 5 4.5 4 3.5 3 (on 2.5 2 1.5 1 0.5 0 -2 0 2 4 6 8 10 B